P2-1 Transaction costs You would like to purchase one Class A share of Berkshire Hath- away through your Scottrade brokerage account. Scottrade charges a $7 commission for online trades. You log into your account, check the real-time quotes for Berk- shire Hathaway (you see a bid price of $262,850 and an ask price of S263,770) and submit your order. a. What is the current bid/ask spread for Berkshire Hathaway Class A shares? b. If Scottrade routes your buy order to the NYSE, where Berkshire Hathaway is listed, whats the potential minimum your total transaction costs will be? c. If, instead, Scottrade routes your buy order to the Nasdag, where Berkshire Hathaway is not listed, whats the potential maximum your total transaction costs will be? d. Regardless of how your trade is executed, based on the bid/ask spread what is the market value of your trade?
Answer #1

a)Bid ask spread= ask price-bid price
=263770-262850=920
b)the minimum transaction cost per share will be the =bid/ask spread+ commission on online
=920+7=927
c)As here the shares not listed in NASDAQ the cost will be only bid/ask spread and it is 920
d)market value of trade=(bid+ask)/2
=(262850+263770)/2=263310